A 0.105m HF (aq) solution shows a freezing point of -0.198 C. Calculate the % dissociation of HF.
1
Expert's answer
2019-09-19T03:58:45-0400

The freezing point depression equation is:


"\Delta T_f= i\times K_f\times m"

The only value we do not know is i:


"0.198 = i\times 1.86 \times 0.105"


"i = 1.014"


% ionozation "= (i-1)\times 100 \%"


"% ionization = (i-1)\times 100%" % ionization = "=(1.014-1)\times 100 \% = 14\%"

"% ionization = (i-1)\times 100%"

The HF is about 14% dissociated.